It has been observed that the mean breaking strength of a particular brand of thread is 9.72N with a standard deviation of 1.4N. When a sample of 30 pieces is taken it showed a mean breaking strength of 8.93N. Can it be concluded that at the 5% level the thread has become inferior?
1
Expert's answer
2020-11-18T19:30:04-0500

The provided sample mean is 

"\bar X = 8.93"  and the known population standard deviation is σ=1.4, and the sample size is n = 30.

The following null and alternative hypotheses need to be tested:

"\\

H_0:μ≥9.72\\

Ha:μ<9.72"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

The z-statistic is computed as follows:

"z=\dfrac{\overline{x}-\mu_0}{\frac{\sigma}{\sqrt{n}}}"


"z=\dfrac{8.93-9.72}{\frac{1.4}{\sqrt{30}}}" = −3.091

Using Z table, are


The p-value is p = 0.001, P(Z<=-3.091)= 0.001,since p=0.001<0.05, it is concluded that the null hypothesis is rejected.

It is concluded that the null hypothesis Ho is rejected. Therefore, there is enough evidence to claim that the population mean μ is less than 9.72, at the 0.05 significance level.